OpenStudy (anonymous):

what is the equation in point-slope form of the line passing through (-2,0) and (2,8)

OpenStudy (anonymous):

First find the slope of the two points. Do you know how?

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

slope(m) = (y2-y1)/(x2-x1)

OpenStudy (jdoe0001):

\(\huge m=\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

omg i cant believe that i didnt know that lol thank you

OpenStudy (anonymous):

just because you find the slope...that is not the answer

OpenStudy (anonymous):

Once you have the slope you can use the formula... y-y1 = m(x-x1)

OpenStudy (anonymous):

That is true kelliegirl33. One thing at a time though.